# OpenDecoder: Open Large Language Model Decoding to Incorporate Document Quality in RAG

### v1

@@ Abstract.
Information Retrieval, Retrieval-Augmented Generation, Robust Question Answer, Decoding Paradigm, Large Language Model
journalyear: 2026copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2026; April 13-17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2023 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesccs: Information systems Information retrievalccs: Computing methodologies Artificial intelligence
@@ 2.1. Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Gao et al., 2023) aims to retrieve external resources to supplement LLMs to generate a response, showing significant advantages in knowledge-intensive tasks (Guu et al., 2020; Kang et al., 2023; Dong et al., 2025; Zhang et al., 2025b). Earlier RAG methods follow the “Retrieve-then-Read” framework (Lewis et al., 2020; Izacard et al., 2023) by adopting a retriever to search for relevant information from external resources based on the user’s query. To further enhance RAG performance, subsequent studies focus on refining retrieval quality through techniques such as query reformulation (Ma et al., 2023; Mo et al., 2023), re-ranking (Sun et al., 2023; Yu et al., 2024; Meng et al., 2026), and noise filtering as intermediate steps (Jin et al., 2024; Qian et al., 2025; Mo et al., 2026), thereby improving the relevance of documents before they are appended to LLMs’ input.
@@ 2.2. Decoding Optimization in LLMs
Prompting (Liu et al., 2023) the advanced LLMs is a simple and effective way to instruct them to generate answers, where the answer decoding highly relies on the designed prompt and internal attention mechanism. Existing literature optimizes the decoding procedure of LLMs on various aspects. For efficiency, Performers (Choromanski et al., 2021) propose compressed attention, reducing attention complexity from quadratic to linear. StreamingLLM (Xiao et al., 2024) leverages attention sinks to decrease Key-Value cache memory for long-context generation. For effectiveness, a series of studies (Zhang et al., 2024; Chan et al., 2024; Yue et al., 2025; Tan et al., 2025; Zhang et al., 2025a) investigate how to leverage inference scaling and deep reasoning for RAG decoding. A recent study REFRAG (Lin et al., 2025) rethinks RAG-based decoding and proposes an optimized architecture to compress only a small subset of retrieved documents that are directly related to the query for effective and efficient decoding. For faithfulness, the existing studies aim to detect and manage misinformation within retrieved documents (Zhou et al., 2024), such as explicitly identifying and resolving knowledge conflicts (Zhang et al., 2025c; Wang et al., 2025a; Deng et al., 2025). These studies focus on selecting relevant and reliable information for LLM input, which still operate in the way that has been trained, by assuming the input information to be relevant. In contrast, in our approach, we modify the attention mechanism according to the relevance of retrieved information. Such an approach has not been proposed in the literature.
@@ 5.5. Noise Tolerance of Input Top-K
5.6. Investigation of Scaling Model Size
We further investigate the impact of scaling up model size for vanilla SFT and our OpenDecoder. The results in the noisy evaluation setting are depicted in Figure 6. Overall, both the SFT and our proposed approaches benefit from larger model sizes, suggesting that larger models are more capable of tolerating contextual noise, which aligns with prior studies (Kaplan et al., 2020). Moreover, the effectiveness of leveraging explicit indicators to influence answer generation becomes more pronounced with larger models, whereas smaller models (e.g., 1.5B) do not consistently achieve better performance across all datasets. These observations indicate that effectively integrating external signals with internal LLM reasoning processes is a non-trivial task that demands higher model capacity. A similar trend is observed when aggregating multiple guidance score features, implying that this aggregation process also requires implicit learning during training. Therefore, designing more sophisticated learning objectives to better incorporate this aggregation mechanism and employing larger backbone models for training OpenDecoder could further enhance performance, which we leave for future work. Results about the evaluation in the other two settings are provided in Appendix C.
@@ Appendix B Baseline Details
Appendix C More Results on Model Scaling
The results in the normal evaluation and extreme noisy setting are depicted in Figure 7 and Figure 8, respectively. Overall, similar trends are observed in the noisy evaluation setting of Sec. 5.6, where larger models are more capable of tolerating contextual noise. Besides, the improvement in scaling model size is more pronounced in complex QA datasets than in general ones, indicating that a larger model may be equipped with a more powerful reasoning ability implicitly.

### v2

@@ Abstract.
Information Retrieval, Retrieval-Augmented Generation, Robust Question Answer, Decoding, Large Language Model
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792524isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Information retrievalccs: Information systems Data management systemsccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning
@@ 2.1. Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Gao et al., 2023) aims to retrieve external resources to supplement LLMs to generate a response, showing significant advantages in knowledge-intensive tasks (Guu et al., 2020; Kang et al., 2023; Dong et al., 2025). Earlier RAG methods follow the “Retrieve-then-Read” framework (Lewis et al., 2020; Izacard et al., 2023) by adopting a retriever to search for relevant information from external resources based on the user’s query. To further enhance RAG performance, subsequent studies focus on refining retrieval quality through techniques such as query reformulation (Ma et al., 2023; Mo et al., 2023), re-ranking (Sun et al., 2023; Yu et al., 2024), and noise filtering as intermediate steps (Jin et al., 2024; Qian et al., 2025), thereby improving the relevance of documents before they are appended to LLMs’ input.
@@ 2.2. Decoding Optimization in LLMs
Prompting (Liu et al., 2023) the advanced LLMs is a simple and effective way to instruct them to generate answers, where the answer decoding highly relies on the designed prompt and internal attention mechanism. Existing literature optimizes the decoding procedure of LLMs on various aspects. For efficiency, Performers (Choromanski et al., 2021) propose compressed attention, reducing attention complexity from quadratic to linear. StreamingLLM (Xiao et al., 2024) leverages attention sinks to decrease Key-Value cache memory for long-context generation. For effectiveness, a series of studies (Chan et al., 2024; Yue et al., 2025; Tan et al., 2025) investigate how to leverage inference scaling and deep reasoning for RAG decoding. A recent study REFRAG (Lin et al., 2025) rethinks RAG-based decoding and proposes an optimized architecture to compress only a small subset of retrieved documents that are directly related to the query for effective and efficient decoding. For faithfulness, the existing studies aim to detect and manage misinformation within retrieved documents (Zhou et al., 2024), such as explicitly identifying and resolving knowledge conflicts (Zhang et al., 2025; Wang et al., 2025a; Deng et al., 2025). These studies focus on selecting relevant and reliable information for LLM input, which still operate in the way that has been trained, by assuming the input information to be relevant. In contrast, in our approach, we modify the attention mechanism according to the relevance of retrieved information. Such an approach has not been proposed in the literature.
@@ 6. Conclusion
Acknowledgements.
This work was partly supported by a discovery grant from the Natural Science and Engineering Research Council of Canada (NSERC) and the Canada Research Chair on natural language information processing and applications.
@@ Appendix B Baseline Details
Appendix C Investigation of Scaling Model Size
We further investigate the impact of scaling up model size for vanilla SFT and our OpenDecoder. The results in the noisy evaluation setting are depicted in Figure 6. Overall, both the SFT and our proposed approaches benefit from larger model sizes, suggesting that larger models are more capable of tolerating contextual noise, which aligns with prior studies (Kaplan et al., 2020). Moreover, the effectiveness of leveraging explicit indicators to influence answer generation becomes more pronounced with larger models, whereas smaller models (e.g., 1.5B) do not consistently achieve better performance across all datasets. These observations indicate that effectively integrating external signals with internal LLM reasoning processes is a non-trivial task that demands higher model capacity. A similar trend is observed when aggregating multiple guidance score features, implying that this aggregation process also requires implicit learning during training.
The results in the normal evaluation and extreme noisy setting are also depicted in Figure 7 and Figure 8, respectively. Overall, similar trends are observed in the noisy evaluation setting, where larger models are more capable of tolerating contextual noise. Besides, the improvement in scaling model size is more pronounced in complex QA datasets than in general ones, indicating that a larger model may be equipped with a more powerful reasoning ability implicitly. Therefore, designing more sophisticated learning objectives to better incorporate this aggregation mechanism and employing larger backbone models for training OpenDecoder could further enhance performance, which we leave for future work.